OpenStudy (anonymous):

For a holiday party, Kathleen is making a six-layer torte, which consists of six 6-inch diameter cake layers with a cream topping on each layer. She knows that one recipe of cream will cover 45 square inches of cake. How many batches of the topping must she make to complete her torte? ( A = p )(Use p = 3.14 as an approximate value) 4 batches 3 batches 16 batches 15 batches

OpenStudy (shane_b):

Do you know how to get the area of a circle?

OpenStudy (anonymous):

I was never good at math. -.-

OpenStudy (shane_b):

No problem. The area of a circle is:\[A_{circle}=\pi * radius^2\]That's the same as \[A_{circle}=\pi (\frac{diameter}{2})^2\]So each cake layer will have an area of \[A=\pi 3^2=28.27in^2\]And the cake has 6 layers so the total area becomes:\[(6)(28.27in^2)=169.62in^2\]How many batches would it take to fill that amount of area if each batch is 45 in^2?

OpenStudy (anonymous):

3.75^2 right? Which is basically 4?

OpenStudy (shane_b):

You got it :)

OpenStudy (anonymous):

can you help me with another real fast? [(1/2)^3 - (1/4)2]2^3

OpenStudy (shane_b):

Ok. To be clear, is that this?\[[(\frac{1}{2})^3-\frac{1}{4}2)]2^3\]

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

Ack Im sorry... the 2 after the 1/4 is an exponent

OpenStudy (anonymous):

Just re-looked at the homework problem

OpenStudy (shane_b):

I kinda thought it would be :) Ok, take it in pieces. Starting inside the brackets... \[(\frac{1}{2})^3=\frac{1^3}{2^3}=\frac{1}{8}\]and\[(\frac{1}{4})^2=\frac{1^2}{4^2}=\frac{1}{16}\]And you should know that 2^3 = 8. So now we have this:\[8(\frac{1}{8}-\frac{1}{16})=?\]All you need to do is distribute the 8 and perform the subtraction.

OpenStudy (anonymous):

1/2

OpenStudy (shane_b):

Yep

OpenStudy (anonymous):

THank you much

OpenStudy (shane_b):

yw